OpenStudy (liljj1421):

what is the slop of y+1/4 x=2

Mehek (mehek14):

move \(\dfrac{1}{4}x\) to the other side first so you would subtract it \(y=-\dfrac{1}{4}x+2\) slope is the number attached to x, which is \(-\dfrac{1}{4}\)

OpenStudy (liljj1421):

how would you graph it

Mehek (mehek14):

plug any value for x and find y

rebeccaxhawaii (rebeccaxhawaii):

solve for y

rebeccaxhawaii (rebeccaxhawaii):

 then m is slope and b is y intercept

OpenStudy (liljj1421):

so what #s do you plug in